repare an amortization schedule for a five-year loan of $45,000. The interest rate is 8% per year, and the loan calls for equal annual payments. (Do not round intermediate calculations. Enter all amount as positive value. Round the final answers to 2 decimal places. Leave no cells blank - be certain to enter "0" wherever required.)

Year Beginning
Balance
Total
Payment
Interest
Payment
Principal
Payment
Ending
Balance
1 $ $ $ $ $
2
3
4
5

How much interest is paid in the third year? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Interest paid           $

How much total interest is paid over the life of the loan? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Total interest           $
Answer #1

Year 1) 1 2 3 4 5 Beginning Total Interest Principal Ending balance Payment payment payment balance (a) (b) (c=a * 8%) (d=b-c